# V Holešovičkách 2, 18000 Praha 8, Czech Republic

### v1

@@ 4.4.2 Lensing regimes in parameter space: low convergence parameter
The four primary boundaries seen in the left panel all extend further beyond its range, intersecting all vertical and horizontal lines, i.e., the corresponding transitions occur for any fixed value of $\kappa_{\text{s}}$ or $e\in(0,1)$. Note that the boundaries do not intersect within the parameter limits of the left panel, so that the sequence of the corresponding transitions remains the same. In order of increasing eccentricity, the lowest bold solid black boundary marks the piercing of the radial caustic by the tangential caustic along the major axis. The equation identifying this condition can be derived from the lens Eq. (49). The piercing occurs when the two caustic points along the major axis each generate one image on the tangential critical curve (on the same side from the center) and another image on the radial critical curve (on the opposite side):
@@
obtained by differentiating Eq. (51). Note that while Eq. (57) depends on the inflection-point coordinate $x_{2}$, $e$, and $\kappa_{\text{s}}$, in Eq. (58) $\kappa_{\text{s}}$ cancels out. It is thus advantageous to choose a value of $e$, solve this equation numerically to obtain $x_{2}$, and use both to directly compute the corresponding value of $\kappa_{\text{s}}$ from Eq. (57).
@@
4.4.3 Lensing regimes in parameter space: intermediate and high convergence parameter
@@
Note that the described “bulge” boundary peaks at $(\kappa_{\text{s}},e)\approx(2.296,0.9881)$, which also identifies the highest-eccentricity parameter combination within the range of Fig. 8 that has an entirely convex tangential critical curve. To the right from the green point in Fig. 8 the dashed thin green boundary describes a different version of the “peanut” transition. When crossed from lower to higher eccentricities it corresponds to the disappearance of two symmetric pairs of inflection points along the minor axis, turning the bulged critical curve into a peanut-shaped one.
@@
Figure 10: High-$\kappa_{\text{s}}$ lensing-regime boundaries of the ellipsoidal NFW lens mapped in its parameter space, as described in Sect. 4.4.3, plotted here for ellipticities $0.651-\sqrt{1-e^{2}}>$ 0.65 and for convergence parameters up to $\kappa_{\text{s}}=140$. Notation same as in Fig. 8, with the additional thin dashed purple boundary marking the minor-axis radial-peanut transition.
@@
The second new boundary, plotted in bold dashed green at the top right corner of the right panel of Fig. 8 indicates two butterfly metamorphoses occurring at the minor-axis cusps of the tangential caustic. Inside the region to the right of the boundary, these cusps are replaced by small three-cusped self-intersecting butterfly structures. An example corresponding to the $(\kappa_{\text{s}},e)\approx(8,0.9986)$ parameter combination marked by the right blue dot is shown in the right column of Fig. 9, with an inset showing a blow-up of the vicinity of the top cusp revealing the butterfly. These caustic tangles lead to very localized changes in the number of images: five images in the butterfly “wings”, and seven images in the butterfly “proboscis” (given that the structure lies inside the oval radial caustic). The described butterfly metamorphosis occurs for convergence-parameter values $\kappa_{\text{s}}\gtrapprox 7.185$. We identified and located this transition by studying the changes in the number of cusps using the image-plane analysis described in Appendix C. Note that while the vertical extent of the butterfly features in the example in Fig. 9 is merely $\approx 0.06\%$ of the vertical extent of the tangential caustic, their prominence grows with increasing convergence parameter so that at $\kappa_{\text{s}}=100$ and higher eccentricities it exceeds $10\%$ of the vertical caustic extent.
The structure of the parameter-space map at higher convergence-parameter values, further to the right of the range of Fig. 8 is shown in Fig. 10 for $\kappa_{\text{s}}\in[0,140]$. We show here only the region of higher ellipticity, $1-\sqrt{1-e^{2}}\geq 0.65$, corresponding to eccentricities $e\gtrsim 0.937$. No boundaries occur at lower eccentricity for $\kappa_{\text{s}}>5$ within the range of Fig. 10. The plot illustrates the changes of the boundaries from Fig. 8 as well as the appearance of another new boundary. The thin dashed purple boundary extending to the right from $\kappa_{\text{s}}\approx 41.6$ corresponds to another “peanut” transition on the radial critical curve. In this one two symmetric pairs of inflection points appear at the minor axis of the mass distribution, so that to the right of the boundary the radial critical curve is pinched in vertically.
Regarding the sequence of the boundaries, a closer inspection reveals there are only two boundaries that gradually intersect others within the range of Fig. 10; the remaining boundaries do not cross each other. The solid black cusp-piercing boundary intersects the green dotted “bulge” boundary at $\kappa_{\text{s}}\approx 7.37$, the bold green dashed butterfly boundary at $\kappa_{\text{s}}\approx 26.1$, and the purple dashed radial “peanut” boundary at $\kappa_{\text{s}}\approx 121$. The solid purple radial “peanut” boundary crosses the thin green dashed “peanut” boundary at $\kappa_{\text{s}}\approx 2.98$, the bold green dashed butterfly boundary at $\kappa_{\text{s}}\approx 7.62$, and the dashed purple radial “peanut” boundary $\kappa_{\text{s}}\approx 78.3$. Over most of the range of Fig. 10 the dotted green “bulge” boundary represents the first transition when increasing the eccentricity from zero. The boundary drops to a minimum at $(\kappa_{\text{s}},e)\approx(62.6,0.950)$, then rises slowly with increasing $\kappa_{\text{s}}$. In general, note the entirely different sequence of nine transitions with increasing eccentricity at the right edge of Fig. 10 in comparison with the generic low-convergence sequence of four transitions as seen in the left panel of Fig. 8.
We haven’t explored the parameter space for higher values of $\kappa_{\text{s}}$ beyond the range of Fig. 10 comprehensively. Overall, the boundaries tend to level out, their rate of change declines with increasing $\kappa_{\text{s}}$. The lower part of the new purple dashed peanut boundary from Fig. 10 declines monotonically asymptotically to $e\approx 0.9945$. The declining lower part of the bold green dashed butterfly boundary eventually crosses the slowly rising dotted green bulge boundary at $(\kappa_{\text{s}},e)\approx(8000,0.956)$. For higher values of $\kappa_{\text{s}}$ the butterfly boundary marks the first transition to occur when increasing the eccentricity from zero. However, at such high values of the convergence parameter these results purely illustrate the changing properties of the simple lens model.
@@
Figure 11: Radii of characteristic curves of the spherical NFW lens, plotted as a function of the convergence parameter up to $\kappa_{\text{s}}=2.1$ (left panel) and up to $\kappa_{\text{s}}=8$ (right panel). Included are the radii of the tangential ($x_{\text{T}}$) and radial ($x_{\text{R}}$) critical curves, of the tangential ($y_{\text{T}}$) and radial ($y_{\text{R}}$) caustics, and of the unit-convergence circle ($x_{0}$). The radius of the radial limiting curve ($x_{\text{R}\infty}\approx 1.32$) is marked along the right edge of the right panel. The five vertical dashed lines mark the values of $\kappa_{\text{s}}$ for which critical-curve and caustic dimensions are plotted for the ellipsoidal NFW lens in Fig. 12 and Fig. 14, respectively.
@@ 4.4.4 Critical-curve and caustic radii in the circularly symmetric case
In Sect. 4.4.2 and Sect. 4.4.3 we described the range of possible shapes of the critical curves and caustics. In this and the following two sections we study the linear dimensions and areas of their tangential and radial components.
For initial orientation we plot in Fig. 11 the radii of the critical curves and caustics in the $e=0$ case for convergence-parameter intervals $\kappa_{\text{s}}\in[0,2.1]$ (left panel) and $\kappa_{\text{s}}\in[0,8]$ (right panel), matching the horizontal extent of the panels in Fig. 8. Note that this case includes purely spherical NFW lenses as well as ellipsoidal NFW lenses that are circular in projection, as discussed in the text following Eq. (7). The plots are extensions of Fig. 3 from Karamazov et al. (2021) to higher values of $\kappa_{\text{s}}$.
@@ 4.4.5 Critical-curve dimensions in the general ellipsoidal case
The same five convergence-parameter values are also marked for convenience in Fig. 11 and Fig. 8 by the thin vertical dashed lines. The former may serve as a reference for the $e=0$ starting points along the left axes of the plots in the left column of Fig. 12. The latter is useful for keeping track of the transitions between different lensing regimes encountered while increasing the ellipticity from zero to one.
@@
With increasing $\kappa_{\text{s}}$ the growth of the radial critical curve slows down as it converges to a radial limiting curve, as mentioned above and described in Appendix C. The horizontal and vertical “radii” of this curve are plotted in Fig. 13 as a function of ellipticity. The radii start at zero ellipticity from the value $x_{\text{R}\infty}\approx 1.32$ marked in the right panel of Fig. 11. Both decrease monotonically with ellipticity: the horizontal reaches $x_{1}=1$; the vertical $x_{2}=0$ at $e=1$. The radial limiting curve shrinks from a circle at $e=0$ to an increasingly horizontally elongated shape, converging at $e=1$ to a horizontal line segment extending one scale radius to either side of the halo center. Note that although the horizontal and vertical radii of radial critical curves converge to the lines plotted in Fig. 13 as $\kappa_{\text{s}}\to\infty$, their horizontal radius always shrinks to zero in the limit $e\to 1$, as mentioned above.
@@ 4.5.1 Magnification and flattening
In the lower three panels of Fig. 15, the shear along the major and minor axes is marked by the red and blue lines, respectively. The green line with the semi-major-axis ticks corresponds to $\theta=\pi/4$, a line dividing the quarter-ellipse into equal-area parts. We see that in the inner region of the halo covering nearly the full extent of the plotted CS diagrams, the shear is lowest along the minor axis and highest along the major axis. In the outer region extending asymptotically to infinity, the order is reversed: minimum shear occurs along the major axis and maximum along the minor axis. In the plotted examples, the transition between the inner and outer regions occurs between $a=1$ and $a=2$. An inspection of the shear plots in Fig. 5 and Fig. 4 confirms these results.
@@
We illustrate this property in Fig. 16, for simplicity on the example of the circularly symmetric case from the top panel of Fig. 15. The five green curves correspond to convergence-parameter values $\kappa_{\text{s}}\in\{0.125,0.25,0.5,1,2\}$ as marked along the right side of the plot (the same values were used in Fig. 12 and Fig. 14). For $\kappa_{\text{s}}=0.125$ all points (including their $a$ values) are moved two times closer to the origin with respect to the $\kappa_{\text{s}}=0.25$ case, for $\kappa_{\text{s}}=0.5$ all points are moved two times further from the origin, etc. For orientation, we connected points with $a=1$ on all the curves by the top dashed orange line, and added two more dashed orange lines similarly for all $a=0.1$ and all $a=0.01$ points.
@@
4.5.2 Orientation
Here we use the term “image orientation” as a shorthand for the orientation of the image distortion, or, even more specifically, for the orientation of the major axis of an elliptical image of a small circular source [^2] Based on its equality to the phase in weak lensing, Karamazov & Heyrovský (2022) use the term “weak phase” for the same quantity.. As such the orientation is primarily defined by the lens phase $\phi$, described in Sect. 4.3 and illustrated in Fig. 4 and Fig. 5. The secondary factor is the convergence: outside the unit-convergence ellipse where $\kappa<1$, the image orientation is equal to the phase; inside the unit-convergence ellipse where $\kappa>1$, the image orientation is perpendicular to the phase. We quantify the orientation by the polar angle in the lens plane, measured counter-clockwise from the major axis of the mass distribution modulo $\pi$, using the corresponding value from the interval $[-\pi/2,\pi/2]$.
@@ 4.5.2 Orientation
The third panel of Fig. 17 illustrates the case when the convergence parameter is equal to the umbilic value of $\kappa_{\text{s}}$ at the given eccentricity. Here the zero-shear points lie on the unit-convergence ellipse and thus they form the umbilic points. Outside the unit-convergence ellipse horizontally oriented images appear along the minor axis, while vertically oriented images appear along the major axis. Along the tangential critical curve image orientation deviates by barely more than ±$\pm\pi/4$ from the vertical, so that while they are oriented tangentially near the major axis, when approaching the minor axis they are oriented nearly perpendicularly to the critical curve. Inside the unit-convergence ellipse horizontally oriented images appear along both the major and the minor axis, vertically oriented images are entirely absent here and the image orientation deviates by less than ±$\pm\pi/4$ from the horizontal. Along the radial critical curve the images are oriented radially near the major axis, and when approaching the minor axis they are oriented nearly perpendicularly to the critical curve. In either half-plane, constant-orientation contours with orientation in the interval $[-\pi/4,\pi/4]$ meet at the umbilic points, and there is no change in orientation when crossing the points. Note that there are no vertically oriented images in the vicinity of the umbilic points, all images deviate by no more than ±$\pm\pi/4$ from the horizontal.
@@ 4.5.3 Configuration examples
In the top left example, the lens parameters $(\kappa_{\text{s}},e)=(0.25,0.9)$ correspond to the fourth CS diagram in Fig. 15 and the second image-orientation map in Fig. 17. The circular sources with radii 0.004 (red) and 0.002 (yellow) are centered at the point $(y_{1},y_{2})=(0,0.0345)$, so that the yellow source lies fully between the radial and tangential caustics while the red source crosses both caustics. Due to its position, the yellow source has three full separate images. The top-most part of the red source lying outside the tangential caustic has one image, the part between the caustics has three, and the lowest part lying within the radial caustic has five images. One full image of the red source lies along the minor axis above the critical curves, while all other images are merged into one large bell-shaped macro-image under the halo center. Note the horizontal orientation of the middle yellow image, which is tangent to the radial critical curve, and the vertical orientation of the bottom yellow image, which is perpendicular to the tangential critical curve. Both are in agreement with the second image-orientation map in Fig. 17, as discussed in Sect. 4.5.2. Similar distortions in the corresponding parts of the red macro-image lead to its peculiar shape.
@@
In the top right example in Fig. 18, the lens parameters $(\kappa_{\text{s}},e)=(0.398,0.9)$ represent the umbilic combination seen also in the third image-orientation map of Fig. 17. The circular sources with radii 0.01 (red) and 0.005 (yellow) are centered at the point $(y_{1},y_{2})=(0,0.09)$, so that the yellow source lies fully inside both caustics while the red source crosses both caustics. The yellow source has five full separate images. The narrow parts of the top edge of the red source lying outside the tangential caustic have one image, the narrow parts between the caustics have three, and the rest of the red source lying inside the radial caustic has five images. One full image of the red source lies along the minor axis above the critical curves, while all other images are merged into one large bumpy ring-like macro-image around the umbilic point (rather than around the halo center). Note the horizontal orientation of the yellow images above and below the umbilic point and the corresponding dominant horizontal distortion of the top and bottom parts of the red ring-like image. The left and right “bumps” on the ring correspond to the abrupt change in orientation along the unit-convergence ellipse, as seen in the third image-orientation map in Fig. 17.
In the bottom left example in Fig. 18, the lens parameters $(\kappa_{\text{s}},e)=(0.5,0.9)$ correspond to the fourth image-orientation map in Fig. 17. The circular sources with radii 0.012 (red) and 0.006 (yellow) are centered at the point $(y_{1},y_{2})=(0,0.14)$, so that the yellow source lies fully inside both caustics while the red source lies inside the radial caustic but crosses the tangential caustic. The yellow source has five full separate images. The narrow parts of the top edge of the red source lying outside the tangential caustic have three images, and the rest of the red source lying inside both caustics has five images. One full image of the red source lies along the minor axis above the critical curves, another full image lies inside the radial critical curve, and all other images are merged into one large boomerang-like macro-image. Note the horizontal orientation of all yellow images along the minor axis, with the lowest one oriented tangentially to the tangential critical curve, and the near-perpendicular orientation of the two side images. The top of the red boomerang-like macro-image is distorted and flattened due to its location inside the unit-convergence ellipse, as seen in the fourth image-orientation map in Fig. 17. In comparison with the top left example, the image configuration is dramatically different even though the critical-curve geometry is superficially similar.

### v2

@@ 4.4.2 Lensing regimes in parameter space at low $\kappa_{\text{s}}$
The four primary boundaries seen in the left panel all extend further beyond its range, intersecting all vertical and horizontal lines, i.e., the corresponding transitions occur for any fixed value of $\kappa_{\text{s}}$ or $e\in(0,1)$. We note that the boundaries do not intersect within the parameter limits of the left panel, so that the sequence of the corresponding transitions remains the same. In order of increasing eccentricity, the lowest bold solid black boundary marks the piercing of the radial caustic by the tangential caustic along the major axis. The equation identifying this condition can be derived from the lens Eq. (49). The piercing occurs when the two caustic points along the major axis each generate one image on the tangential critical curve (on the same side from the center) and another image on the radial critical curve (on the opposite side):
@@
obtained by differentiating Eq. (51). We note that while Eq. (57) depends on the inflection-point coordinate $x_{2}$, $e$, and $\kappa_{\text{s}}$, in Eq. (58) $\kappa_{\text{s}}$ cancels out. It is thus advantageous to choose a value of $e$, solve this equation numerically to obtain $x_{2}$, and use both to directly compute the corresponding value of $\kappa_{\text{s}}$ from Eq. (57).
@@
4.4.3 Lensing regimes in parameter space at intermediate $\kappa_{\text{s}}$
@@
We note that the described “bulge” boundary peaks at $(\kappa_{\text{s}},e)\approx(2.296,0.9881)$, which also identifies the highest-eccentricity parameter combination within the range of Fig. 8 that has an entirely convex tangential critical curve. To the right from the green point in Fig. 8 the dashed thin green boundary describes a different version of the “peanut” transition. When crossed from lower to higher eccentricities it corresponds to the disappearance of two symmetric pairs of inflection points along the minor axis, turning the bulged critical curve into a peanut-shaped one.
@@
The second new boundary, plotted in bold dashed green at the top right corner of the right panel of Fig. 8 indicates two butterfly metamorphoses occurring at the minor-axis cusps of the tangential caustic. Inside the region to the right of the boundary, these cusps are replaced by small three-cusped self-intersecting butterfly structures. An example corresponding to the $(\kappa_{\text{s}},e)\approx(8,0.9986)$ parameter combination marked by the right blue dot is shown in the right column of Fig. 9, with an inset showing a blow-up of the vicinity of the top cusp revealing the butterfly. These caustic tangles lead to very localized changes in the number of images: five images in the butterfly “wings”, and seven images in the butterfly “proboscis” (given that the structure lies inside the oval radial caustic). The described butterfly metamorphosis occurs for convergence-parameter values $\kappa_{\text{s}}\gtrapprox 7.185$. We identified and located this transition by studying the changes in the number of cusps using the image-plane analysis described in Appendix C. We note that while the vertical extent of the butterfly features in the example in Fig. 9 is merely ${\approx}0.06\%$ of the vertical extent of the tangential caustic, their prominence grows with increasing convergence parameter so that at $\kappa_{\text{s}}=100$ and higher eccentricities it exceeds $10\%$ of the vertical caustic extent.
The shape of the boundaries near $\kappa_{\text{s}}=8$ in the right panel of Fig. 8 indicates that further changes in the structure of the parameter space may be expected at higher values of the convergence parameter. We describe the properties of the ellipsoidal NFW lens at high $\kappa_{\text{s}}$ in Appendix D and discuss the astrophysical relevance of its high-eccentricity features in Section 5.1.
@@
Figure 10: Radii of characteristic curves of the spherical NFW lens, plotted as a function of the convergence parameter up to $\kappa_{\text{s}}=2.1$ (left panel) and up to $\kappa_{\text{s}}=8$ (right panel). Included are the radii of the tangential ($x_{\text{T}}$) and radial ($x_{\text{R}}$) critical curves, of the tangential ($y_{\text{T}}$) and radial ($y_{\text{R}}$) caustics, and of the unit-convergence circle ($x_{0}$). The radius of the radial limiting curve ($x_{\text{R}\infty}\approx 1.32$) is marked along the right edge of the right panel. The five vertical dashed lines mark the values of $\kappa_{\text{s}}$ for which critical-curve and caustic dimensions are plotted for the ellipsoidal NFW lens in Figs. 11 and 13, respectively.
@@ 4.4.4 Critical-curve and caustic radii in the circularly symmetric case
In Sects. 4.4.2, 4.4.3, and Appendix D we described the range of possible shapes of the critical curves and caustics. In this and the following two sections we study the linear dimensions and areas of their tangential and radial components.
For initial orientation we plot in Fig. 10 the radii of the critical curves and caustics in the $e=0$ case for convergence-parameter intervals $\kappa_{\text{s}}\in[0,2.1]$ (left panel) and $\kappa_{\text{s}}\in[0,8]$ (right panel), matching the horizontal extent of the panels in Fig. 8. We note that this case includes purely spherical NFW lenses as well as ellipsoidal NFW lenses that are circular in projection, as discussed in the text following Eq. (7). The plots are extensions of Fig. 3 from Karamazov et al. (2021) to higher values of $\kappa_{\text{s}}$.
@@ 4.4.5 Critical-curve dimensions in the general ellipsoidal case
The same five convergence-parameter values are also marked for convenience in Figs. 10 and 8 by the thin vertical dashed lines. The former may serve as a reference for the $e=0$ starting points along the left axes of the plots in the left column of Fig. 11. The latter is useful for keeping track of the transitions between different lensing regimes encountered while increasing the ellipticity from zero to one.
@@
With increasing $\kappa_{\text{s}}$ the growth of the radial critical curve slows down as it converges to a radial limiting curve, as mentioned above and described in Appendix C. The horizontal and vertical “radii” of this curve are plotted in Fig. 12 as a function of ellipticity. The radii start at zero ellipticity from the value $x_{\text{R}\infty}\approx 1.32$ marked in the right panel of Fig. 10. Both decrease monotonically with ellipticity: the horizontal reaches $x_{1}=1$; the vertical $x_{2}=0$ at $e=1$. The radial limiting curve shrinks from a circle at $e=0$ to an increasingly horizontally elongated shape, converging at $e=1$ to a horizontal line segment extending one scale radius to either side of the halo center. We note that although the horizontal and vertical radii of radial critical curves converge to the lines plotted in Fig. 12 as $\kappa_{\text{s}}\to\infty$, their horizontal radius always shrinks to zero in the limit $e\to 1$, as mentioned above.
@@ 4.5.1 Magnification and flattening
In the lower three panels of Fig. 14, the shear along the major and minor axes is marked by the red and blue lines, respectively. The green line with the semi-major-axis ticks corresponds to $\theta=\pi/4$, a line dividing the quarter-ellipse into equal-area parts. We see that in the inner region of the halo covering nearly the full extent of the plotted CS diagrams, the shear is lowest along the minor axis and highest along the major axis. In the outer region extending asymptotically to infinity, the order is reversed: minimum shear occurs along the major axis and maximum along the minor axis. In the plotted examples, the transition between the inner and outer regions occurs between $a=1$ and $a=2$. An inspection of the shear plots in Figs. 5 and 4 confirms these results.
@@
We illustrate this property in Fig. 15, for simplicity on the example of the circularly symmetric case from the top panel of Fig. 14. The five green curves correspond to convergence-parameter values $\kappa_{\text{s}}\in\{0.125,0.25,0.5,1,2\}$ as marked along the right side of the plot (the same values were used in Figs. 11 and 13). For $\kappa_{\text{s}}=0.125$ all points (including their $a$ values) are moved two times closer to the origin with respect to the $\kappa_{\text{s}}=0.25$ case, for $\kappa_{\text{s}}=0.5$ all points are moved two times further from the origin, etc. For orientation, we connected points with $a=1$ on all the curves by the top dashed orange line, and added two more dashed orange lines similarly for all $a=0.1$ and all $a=0.01$ points.
@@
4.5.2 Orientation
Here we use the term “image orientation” as a shorthand for the orientation of the image distortion, or, even more specifically, for the orientation of the major axis of an elliptical image of a small circular source [^2] Based on its equality to the phase in weak lensing, Karamazov & Heyrovský (2022) use the term “weak phase” for the same quantity.. As such the orientation is primarily defined by the lens phase $\phi$, described in Sect. 4.3 and illustrated in Figs. 4 and 5. The secondary factor is the convergence: outside the unit-convergence ellipse where $\kappa<1$, the image orientation is equal to the phase; inside the unit-convergence ellipse where $\kappa>1$, the image orientation is perpendicular to the phase. We quantify the orientation by the polar angle in the lens plane, measured counter-clockwise from the major axis of the mass distribution modulo $\pi$, using the corresponding value from the interval $[-\pi/2,\pi/2]$.
@@ 4.5.2 Orientation
The third panel of Fig. 16 illustrates the case when the convergence parameter is equal to the umbilic value of $\kappa_{\text{s}}$ at the given eccentricity. Here the zero-shear points lie on the unit-convergence ellipse and thus they form the umbilic points. Outside the unit-convergence ellipse horizontally oriented images appear along the minor axis, while vertically oriented images appear along the major axis. Along the tangential critical curve image orientation deviates by barely more than ±$\pm\pi/4$ from the vertical, so that while they are oriented tangentially near the major axis, when approaching the minor axis they are oriented nearly perpendicularly to the critical curve. Inside the unit-convergence ellipse horizontally oriented images appear along both the major and the minor axis, vertically oriented images are entirely absent here and the image orientation deviates by less than ±$\pm\pi/4$ from the horizontal. Along the radial critical curve the images are oriented radially near the major axis, and when approaching the minor axis they are oriented nearly perpendicularly to the critical curve. In either half-plane, constant-orientation contours with orientation in the interval $[-\pi/4,\pi/4]$ meet at the umbilic points, and there is no change in orientation when crossing the points. We note that there are no vertically oriented images in the vicinity of the umbilic points, all images deviate by no more than ±$\pm\pi/4$ from the horizontal.
@@ 4.5.3 Configuration examples
In the top left example, the lens parameters $(\kappa_{\text{s}},e)=(0.25,0.9)$ correspond to the fourth CS diagram in Fig. 14 and the second image-orientation map in Fig. 16. The circular sources with radii 0.004 (red) and 0.002 (yellow) are centered at the point $(y_{1},y_{2})=(0,0.0345)$, so that the yellow source lies fully between the radial and tangential caustics while the red source crosses both caustics. Due to its position, the yellow source has three full separate images. The top-most part of the red source lying outside the tangential caustic has one image, the part between the caustics has three, and the lowest part lying within the radial caustic has five images. One full image of the red source lies along the minor axis above the critical curves, while all other images are merged into one large bell-shaped macro-image under the halo center. Worth noticing is the horizontal orientation of the middle yellow image, which is tangent to the radial critical curve, and the vertical orientation of the bottom yellow image, which is perpendicular to the tangential critical curve. Both are in agreement with the second image-orientation map in Fig. 16, as discussed in Sect. 4.5.2. Similar distortions in the corresponding parts of the red macro-image lead to its peculiar shape.
@@
In the top right example in Fig. 17, the lens parameters $(\kappa_{\text{s}},e)=(0.398,0.9)$ represent the umbilic combination seen also in the third image-orientation map of Fig. 16. The circular sources with radii 0.01 (red) and 0.005 (yellow) are centered at the point $(y_{1},y_{2})=(0,0.09)$, so that the yellow source lies fully inside both caustics while the red source crosses both caustics. The yellow source has five full separate images. The narrow parts of the top edge of the red source lying outside the tangential caustic have one image, the narrow parts between the caustics have three, and the rest of the red source lying inside the radial caustic has five images. One full image of the red source lies along the minor axis above the critical curves, while all other images are merged into one large bumpy ring-like macro-image around the umbilic point (rather than around the halo center). Worth noticing is the horizontal orientation of the yellow images above and below the umbilic point and the corresponding dominant horizontal distortion of the top and bottom parts of the red ring-like image. The left and right “bumps” on the ring correspond to the abrupt change in orientation along the unit-convergence ellipse, as seen in the third image-orientation map in Fig. 16.
In the bottom left example in Fig. 17, the lens parameters $(\kappa_{\text{s}},e)=(0.5,0.9)$ correspond to the fourth image-orientation map in Fig. 16. The circular sources with radii 0.012 (red) and 0.006 (yellow) are centered at the point $(y_{1},y_{2})=(0,0.14)$, so that the yellow source lies fully inside both caustics while the red source lies inside the radial caustic but crosses the tangential caustic. The yellow source has five full separate images. The narrow parts of the top edge of the red source lying outside the tangential caustic have three images, and the rest of the red source lying inside both caustics has five images. One full image of the red source lies along the minor axis above the critical curves, another full image lies inside the radial critical curve, and all other images are merged into one large boomerang-like macro-image. Worth noticing is the horizontal orientation of all yellow images along the minor axis, with the lowest one oriented tangentially to the tangential critical curve, and the near-perpendicular orientation of the two side images. The top of the red boomerang-like macro-image is distorted and flattened due to its location inside the unit-convergence ellipse, as seen in the fourth image-orientation map in Fig. 16. In comparison with the top left example, the image configuration is dramatically different even though the critical-curve geometry is superficially similar.